OpenStudy (anonymous):

what is a equivalent fraction for 27 over 243?

OpenStudy (anonymous):

27/243 = 3^3 / 3^5 = 1/3^2 = 1/81

OpenStudy (anonymous):

The answer is 1/9

OpenStudy (anonymous):

yes I am sorry 1/3^2 = 1/9

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

1/9

jhonyy9 (jhonyy9):

27/243 = 1/9